Tamara and will sold papayas at the fair. Tamara sold hers for $5 each, will sold his for $7 each. Tamara sold 3 more than will. Together they made $147. How many papayas did they sell altogether?
Answer by stanbon(75887)

Tamara and will sold papayas at the fair. Tamara sold hers for $5 each, will sold his for $7 each. Tamara sold 3 more than will. Together they made $147. How many papayas did they sell altogether?
-----------------
Value Equation:: 5t + 7w = 147 dollars
Quantity Equation:: t = w+3
-------
Substitute for "t" and solve for "w":
-----
5(w+3) + 7w = 147
12w + 15 = 147
12w = 132
w = 11 ( # Will sold)
t = w+3 = 14 (# Tamara sold)
--------------------
Cheers,
Stan H.
===============